OpenStudy (anonymous):

Phyllis invested $9,500, a portion earning a simple interest rate of 1 1/2% per year and the rest earning a rate of 1% per year. After one year the total interest earned on these investments was $135.00. How much money did she invest at each rate?

OpenStudy (cj49):

let the amt invested at .015 be x x*0.015 + (9500-x)*.01=135

OpenStudy (anonymous):

.015*a + .01*b = 135 a + b = 9500

OpenStudy (anonymous):

System of equations, just do what you normally do with them and solve for one variable first, then substitute it back into either equation and solve for the other.

OpenStudy (cj49):

.015x+95-0.1x=135 0.005x=135-95 0.005x=40 x=40/0.005 x=8000 invested @ 1.5% 9500-8000=1500 invested @1%

OpenStudy (anonymous):

.015a + .01b = 135 .01a + .01b = 95 .005a = 40 a = 8000 Plug that back into a + b = 9500 You get b = 1500

OpenStudy (anonymous):

Where a is the 1.5% interest, b is the 1% interest

OpenStudy (anonymous):

But cj49 is correct too with his method

OpenStudy (anonymous):

I just prefer Linear Algebra methods

OpenStudy (anonymous):

@LizzyMcWire ok thank you but i dont get how you got .01a +.01b=95

OpenStudy (anonymous):

or how did you get that?? @cj49

OpenStudy (anonymous):

I multiplied the whole bottom equation by .01

OpenStudy (cj49):

1/100=0.01 1.5/100=0.015